OpenStudy (anonymous):

Got an odd lil question... can I get some help? ~> Suppose you have a scientific calculator where some of the buttons are broken. In fact the only working keys are the digits [0] to [9], [+],[-],[x],[/] and [e^x] and [ln]. Explain how you can use this calculator to calculate the square root of any whole number.

OpenStudy (anonymous):

some numbers ln divide some number = e^x

ganeshie8 (ganeshie8):

yeah but we dont have ln

OpenStudy (anonymous):

Oh crap

jimthompson5910 (jim_thompson5910):

this page explains the Babylonian method. It's basically making an educated guess and fine tuning that guess to get more accurate results http://www.deltacollege.edu/dept/basicmath/Babylonian.htm

OpenStudy (anonymous):

Ooh nice Jim.

OpenStudy (anonymous):

comeonn guyss...

OpenStudy (anonymous):

and you call yourselves professors of mathematics

OpenStudy (anonymous):

:p

jimthompson5910 (jim_thompson5910):

that page has all the answers you need

OpenStudy (anonymous):

it was something to do with using indirectly using ln

OpenStudy (anonymous):

and it wasn't an estimate :/

jimthompson5910 (jim_thompson5910):

on that page, they have the example of sqrt(5) a good guess is 2. Divide 5 by 2 to get 5/2 = 2.5 then notice how 2*2 = 4 is too small and 2.5*2.5 = 6.25 is too big. So the value of sqrt(5) is between 2 and 2.5 average the two: (2+2.5)/2 = 2.25 then you repeat the process all over again with the new guess of 2.25

OpenStudy (anonymous):

i "guess" that works :p

jimthompson5910 (jim_thompson5910):

well I guess you can say \[\Large \sqrt{x} = x^{1/2}\] but I don't know how to tie it in with \(\Large e^x\) or natural logs. So I'd just stick with the Babylonian method.

OpenStudy (anonymous):

how about if you had [ln]?

OpenStudy (anonymous):

it's okay guys don't fry your brains :p

jimthompson5910 (jim_thompson5910):

still don't know, but I'll think it over

OpenStudy (anonymous):

should be easy if you have [ln]

OpenStudy (anonymous):

okay thanks anyways <3

OpenStudy (anonymous):

oh noes

OpenStudy (anonymous):

I believe in you :)

jimthompson5910 (jim_thompson5910):

ok if you had both e^x and ln(x) buttons, then you can do this \[\Large \sqrt{x} = e^{\ln(\sqrt{x})} \] \[\Large \sqrt{x} = e^{\ln(x^{1/2})} \] \[\Large \sqrt{x} = e^{\frac{1}{2}\ln(x)} \] ------------------------------------------------------- So for example, \[\Large \sqrt{x} = e^{\frac{1}{2}\ln(x)} \] \[\Large \sqrt{5} = e^{\frac{1}{2}\ln(5)} \] \[\Large \sqrt{5} \approx 2.23606797750042 \]

OpenStudy (anonymous):

now wasn't that simple :p